# On Diverse Preferences for Large Language Model Alignment

### 12 Dec 2023 v1

@@ 3.1 Experiment Setup
Datasets. We utilize open human preference datasets for reward modeling, which include Helpful&Harmless (Bai et al., 2022a), OASST1 (Köpf et al., 2023), Webgpt (Nakano et al., 2021a), and Summarize (Stiennon et al., 2020). To avoid the preference unbalance from the size of training sets, we randomly sample 17,106 samples from Helpful, Harmless, OASST1 and Summarize training sets to align with the Webgpt dataset. Besides, we note the unbalance of training data can be addressed by resampling methods (Estabrooks et al., 2004). The corresponding datasets contain (2,352, 2,312, 957, 901, 3,620) test samples respectively.
Baselines. We compare our method with conventional fine-tuning strategies (Dong et al., 2023) in the training language models, which include sequential training and hybrid training. We note that hybrid training policy is essentially a special case of MORE with a static $\lambda=\frac{1}{K}$. Moreover, the sequential training policy only samples a batch of data from single preference datasets during training. We also compare with the naive ensemble, which merely averages the reward values of independently trained rewards following equation 5.
Evaluation metric. We use the preference accuracy on test datasets for each domain. If an RM $r(x,y_{w})>r(x,y_{l})$ for a test sample $(x,y_{w},y_{l})$, we denote it as a correct prediction. The preference accuracy is then computed as the proportion of correct predictions within all testing response pairs. However, preference accuracy only provides pairwise comparisons of responses and does not reflect the degree of preference for each response. Following Bai et al. (2022b); Cheng et al. (2023b), we examine the probability calibration to test if the learned RMs accurately represent the human preference distribution. This is measured by the Expected Calibration Error (ECE) (Naeini et al., 2015; Zhu et al., 2023).
Training setup. Due to the limitation of computational resources, we only select Pythia-1.4B (Biderman et al., 2023) as the LM base for RM training. Experiments with a large scale of the base model are in progress. We use the last token embedding of the output hidden states as the pooled hidden representation, then add one linear layer (RM head) with the scale-value output on it to predict reward scores. All RM training batch size is set to 64. The max input sequence length is 512. All RMs are fine-tuned with one epoch. We use optimizer AdamW with learning rate $5e^{-5}$. We set $\alpha=0.5$ for MORE.
@@ 3.2 Reward Modeling on Multiple Preference Datasets
Main performance. The MORE achieves better testing accuracy and lower ECE values, indicating that the policy provides an accurate reward value on the test. In the last four rows of the table, we train an RM on mixed datasets with different policies. The ensemble, hybrid, and sequential policies show similar test accuracy on single domains and average accuracy. However, the Harmless dataset, which has different preferences compared to the other four datasets, experiences the largest accuracy degeneration with these policies. In contrast, MORE exhibits the lowest performance degeneration on the Harmless test.
Agreement & disagreement analysis. The agreement or disagreement of preferences can be reflected in the testing accuracy. These findings highlight the importance of identifying and understanding the agreement or disagreement across collected datasets in practical applications. By comparing the experimental results of the first six rows, we can observe that when the model is trained on datasets from other domains, the test accuracy on the Harmless dataset decreases. This indicates a disagreement in preferences between the Harmless dataset and the other datasets.
@@
Sequential training analysis. We also find that sequential training is sensitive to the sequence of preference datasets, as shown in Table 2. When adjusting the position of the Harmless dataset while maintaining a fixed sequence for the other datasets (Helpful, Oasst1, Webgpt, Summarize), we observe that sequential training slightly outperforms MORE in terms of testing accuracy when the Harmless dataset is placed at the end. However, MORE performs better on the ECE metric. Therefore, while sequential training has the potential to achieve comparable performance with MORE, it requires additional tuning efforts to adjust the dataset sequence, which can be troublesome when dealing with large and diverse datasets.
Why sequential training works better when Harmless at the last? The training performance is influenced by the proportion of agreement and disagreement data. Previous literature (Touvron et al., 2023a; Bai et al., 2022c) has shown that there is a disagreement between the Helpful and Harmless datasets. Additionally, our observations indicate that the Oasst1, Webgpt, and Summarize datasets share distributional similarities with Helpful (Figure 1) and exhibit similar testing accuracy (Table 1). As a result, the Harmless dataset disagrees with the other datasets, and its proportion in the mixed datasets is relatively small. Due to the smaller proportion of Harmless data, the model updating process allocates less attention to this dataset compared to the others. Sequential training can achieve better performance by adjusting the sequence of datasets. For instance, the recently seen samples are less likely to be forgotten (Wang et al., 2023b). In contrast, MORE successfully enhances performance by adaptively re-weighting losses of different preferences during training using Equation 6. MORE automatically captures the agreement and disagreement through Equation 9 and balances the model updating process accordingly.
3.3 Analysis on Diverse Distribution of Reward Values
@@
Experiment setup. We first partition the testset into an “agreement testset” and a “disagreement testset”. In detail, we let $\{\theta_{\text{helpful}},\theta_{\text{harmless}},\theta_{\text{oasst}},\theta_{\text{webgpt}},\theta_{\text{summ.}}\}$ represent a preference model set. We utilize each preference model in the set to infer every data sample from the testset. We obtain the agreement testset which comprises data samples that agreed by the preference model set following Definition 1. The remaining samples form the disagreement testset. We analyze statistics and distribution of the reward difference (i.e., $r_{\theta}(x,y_{w})-r_{\theta}(x,y_{l})$) in reward values. It is worth noting that a positive reward difference $r_{\theta}(x,y_{w})-r_{\theta}(x,y_{l})>0$ indicates that the model accurately assigns a higher reward value to the winning sample.
Analysis of reward differences on each preference. Our findings indicate that the reward value distributions of different reward models are generally distinct. When examining the RM trained on different single preference datasets, we observe a significant diversity in reward values. In Figure LABEL:fig:reward_dist, we present the reward difference values of the first 50 test samples from the Harmless dataset as an example (Figure LABEL:fig:reward_a). Notably, even on the agreed samples, the baseline model provides different rewards. This observation suggests that the preference distribution learned from a single dataset by the reward model is highly diverse compared to others. Furthermore, Figure LABEL:fig:reward_b provides statistics of the reward difference across the entire test set, further supporting our statements regarding the diversity of reward values.
@@
Analysis of reward difference on training policies. MORE prevents the RM from outputting extreme reward values and maintains the reward accuracy (shown in Table 1). Specifically, we observe the reward values provided by RMs trained with different policies, as depicted in Figure 3. Compared with the sequential training policy, MORE outputs reward values and variance with lower mean and variance. Connecting with our methodology, MORE mitigated the reward drifting. Hence, it tends to output a lower absolute reward signal as more accurate reward values according to ECE in Table 1. Furthermore, the hybrid training policy also yields lower reward values compared to MORE. However, the reward accuracy is poorer in this case, which indicates that compressing the mean and variance of reward differences may degenerate the performance. These results suggest that the too-high or low reward differences may not be good in testing metrics. In contrast, MORE generates appropriate reward values, thereby enhancing the quality of reward modeling.
@@ 4 Discussion
Limitations. In this version, we have exclusively evaluated the MORE approach on the base model Pythia-1.4B. Further works on larger or smaller models are in progress. Besides, we randomly sample training samples from part of the dataset to avoid the effects of unbalanced datasets. The experiment of MORE handling the unbalancing mixed datasets will be presented in the next version. Throughout our experiments, we have found it necessary to utilize equation 8 in MORE, where we set $\alpha=0.5$ for all experiments. Further sensitive study on the value of setting $\alpha$ is in progress.
Future work. We have demonstrated the alignment ability of the reward model using the ECE metric (Zhu et al., 2023). However, the ultimate performance of the LLM, fine-tuned by the RM, remains a crucial aspect of our investigation. In our future work, we plan to complete the RL component in the RLHF pipeline. Additionally, we aim to address several key questions regarding the reward model and MORE: 1) How do the number of diverse datasets and the scaling of base models impact the performance of MORE? 2) Does the RM trained by MORE yield improved results in RL tasks? 3) Can learning feedback from mixed diverse preference datasets enhance the performance of the LLM? Furthermore, we will delve deeper into understanding the implications of preference diversity in reward modeling. Gaining insights into the effects of preference diversity is crucial for designing algorithms that exhibit enhanced data efficiency.
Extension and Applications. The concept of learning shared preferences from mixed diverse preference datasets can be extended to DPO by incorporating the MORE loss (Rafailov et al., 2023). Furthermore, the approach of learning shared preferences through MORE can be implemented as a preference modeling pre-trained (PMP) paradigm (Askell et al., 2021). Subsequently, the RM trained using MORE can be easily fine-tuned to specific preferences (Cheng et al., 2023a). This flexibility allows for the adaptation of our approach to various applications and opens up new possibilities for preference modeling.
@@ 5 Related Work
RLHF framework has become the most popular techniques (Leike et al., 2018; Nakano et al., 2021b; Ouyang et al., 2022; Bai et al., 2022b) to align language models towards helpfulness, and harmlessness. They all utilize an RM to align machine learning systems with human performance, which directly decides the performance of preference alignment. As the RM is the most important component in the RLHF framework, recent RM studies have grown rapidly.
The original goal of RM is to provide a scalar score to a model response and indicate the quality in equation 2, especially helpfulness and harmlessness. Due to the trade-off in quality aspects (Touvron et al., 2023a; Bai et al., 2022b), it can be challenging for a single RM to perform well in all aspects. Our work related to previous works handling multiple rewards and potential disagreement in preferences. For instance, Llama-2 (Touvron et al., 2023a) utilizes two separate RMs, one optimized for helpfulness and another for harmlessness. They mitigate the magnitude bias of the reward scalar with a margin loss, which provides a large margin for pairs with distinct responses, and a smaller one for those with similar responses. Multiple RMs can be utilized as majority voting or averaging (Jaques et al., 2020; Jang et al., 2023) in the PPO (Schulman et al., 2017). Wang et al. (2023a) introduces a Bayesian-based approach called d-PM to align language model with human preferences with disagreement. Cheng et al. (2023a) proposes to train a customized RM from the general RM to avoid disagreement from different preference domains. Furthermore, our theoretical intuition follows recent work DPO (Rafailov et al., 2023) and SLiC-HF (Zhao et al., 2023b) for preference alignment, which explores more straightforward methods to align language models with human preferences. Beyond the methodology, they have shown the RLHF framework is working as likelihood calibration tasks (Deng et al., 2020; Wang et al., 2023a; Azar et al., 2023), which proves that the reward values provided by the RM are also important.
@@ 6 Conclusion
This research paper delves into the problem of reward modeling within the context of mixed diverse preferences datasets. Our experimental findings indicate that RMs trained on different preference datasets exhibit substantial diversity in their reward distributions. This diversity stems from the inherent characteristics of the datasets, posing a challenge for the RM to establish a consensus or accurately represent a shared preference. To tackle this challenge, we propose a straightforward yet highly effective training policy called MORE. The MORE policy is specifically designed to enhance the reward modeling process on diverse preference datasets, to improve the model’s ability to capture and represent the underlying preferences accurately. Through experiments conducted on five diverse datasets, we demonstrate the practical challenges associated with reward modeling. Furthermore, our proposed method outperforms conventional training policies across three different evaluation metrics, highlighting its effectiveness and superiority.
Appendix A Appendix
A.1 Decomposition of ranking loss
@@
A.2 Experiments on Two Diverse Preference Datasets
@@
The results show that MORE implements comparable results with baseline methods in reward accuracy. However, the MORE is not the best on the ECE metric. This is because the number of diverse datasets is small (K=2), resulting in less diverse preference information. Hence, the reward drift mitigation task is solved inefficiently, making MORE degenerate to the vanilla method with rank loss as we demonstrated in Section A.1. Furthermore, we have achieved better results as shown in the main paper with $K=5$, indicating that MORE is efficient in handling diverse preferences in reward modeling.
A.3 Diversity of Reward Values
We provide missing experimental results as supplements of Figure 1.

### 25 Dec 2023 v2

@@ 5.1 Experiment Setup
Datasets
The human preference datasets we used include Helpful&Harmless (Bai et al., 2022a), OASST1 (Köpf et al., 2023), Webgpt (Nakano et al., 2021a), and Summarize (Stiennon et al., 2020). To avoid the preference unbalance from the size of training sets, we randomly sample 17,106 samples from Helpful, Harmless, OASST1 and Summarize training sets to align with the Webgpt dataset. Besides, we note the unbalance of training data can be addressed by resampling methods (Estabrooks et al., 2004). The corresponding datasets contain (2,352, 2,312, 957, 901, 3,620) test samples respectively.
Baselines
We compare our method with conventional fine-tuning strategies (Dong et al., 2023a) in the training language models, which include sequential training and hybrid training. We note that hybrid training scheme is essentially a special case of MORE with a static $\lambda=\frac{1}{K}$. Moreover, the sequential training scheme only samples a batch of data from single preference datasets during training. We also compare with the naive ensemble, which merely averages the reward values of independently trained rewards following equation 5.
Evaluation metric
We use the preference accuracy on test datasets for each domain. If an RM $r(x,y_{w})>r(x,y_{l})$ for a test sample $(x,y_{w},y_{l})$, we denote it as a correct prediction. The preference accuracy is then computed as the proportion of correct predictions within all testing response pairs. However, preference accuracy only provides pairwise comparisons of responses and does not reflect the degree of preference for each response. Following Bai et al. (2022a); Cheng et al. (2023b), we examine the probability calibration to test if the learned RMs accurately represent the human preference distribution. This is measured by the Expected Calibration Error (ECE) (Naeini et al., 2015; Zhu et al., 2023).
Training setup
Due to the limitation of computational resources, we only select Pythia-1.4B (Biderman et al., 2023) as the LM base for RM training. Experiments with a large scale of the base model are in progress. We use the last token embedding of the output hidden states as the pooled hidden representation, then add one linear layer (RM head) with the scale-value output on it to predict reward scores. All RM training batch size is set to 64. The max input sequence length is 512. All RMs are fine-tuned with one epoch. We use optimizer AdamW (Loshchilov & Hutter, 2017) with learning rate $5e^{-5}$. We set $\alpha=0.5$ for MORE.
@@ 5.2 Reward Modeling on Multiple Preference Datasets
Main performance
The MORE achieves better testing accuracy and lower ECE values, indicating that the policy provides an accurate reward value on the test. In the last four rows of the table, we train an RM on mixed datasets with different policies. The ensemble, hybrid, and sequential policies show similar test accuracy on single domains and average accuracy. However, the Harmless dataset, which has different preferences compared to the other four datasets, experiences the largest accuracy degeneration with these policies. In contrast, MORE exhibits the lowest performance degeneration on the Harmless test.
Agreement & disagreement analysis
The agreement or disagreement of preferences can be reflected in the testing accuracy. These findings highlight the importance of identifying and understanding the agreement or disagreement across collected datasets in practical applications. By comparing the experimental results of the first six rows, we can observe that when the model is trained on datasets from other domains, the test accuracy on the Harmless dataset decreases. This indicates a disagreement in preferences between the Harmless dataset and the other datasets.
@@
Sequential training analysis
We also find that sequential training is sensitive to the sequence of preference datasets, as shown in Table 2. When adjusting the position of the Harmless dataset while maintaining a fixed sequence for the other datasets (Helpful, Oasst1, Webgpt, Summarize), we observe that sequential training slightly outperforms MORE in terms of testing accuracy when the Harmless dataset is placed at the end. However, MORE performs better on the ECE metric. Therefore, while sequential training has the potential to achieve comparable performance with MORE, it requires additional tuning efforts to adjust the dataset sequence, which can be troublesome when dealing with large and diverse datasets.
Why sequential training works better when Harmless at the last?
The training performance is influenced by the proportion of agreement and disagreement data. Previous literature (Touvron et al., 2023a; Bai et al., 2022b) has shown that there is a disagreement between the Helpful and Harmless datasets. Additionally, our observations indicate that the Oasst1, Webgpt, and Summarize datasets share distributional similarities with Helpful (Figure 3) and exhibit similar testing accuracy (Table 1). As a result, the Harmless dataset disagrees with the other datasets, and its proportion in the mixed datasets is relatively small. Due to the smaller proportion of Harmless data, the model updating process allocates less attention to this dataset compared to the others. Sequential training can achieve better performance by adjusting the sequence of datasets. For instance, the recently seen samples are less likely to be forgotten (Wang et al., 2023b). In contrast, MORE successfully enhances performance by adaptively re-weighting losses of different preferences during training using Equation 6. MORE automatically captures the agreement and disagreement through Equation 9 and balances the model updating process accordingly.
5.3 Analysis on Diversified Reward Distributions
@@
Experiment setup
We first partition the testset into an “agreement testset” and a “disagreement testset”. In detail, we let $\{\theta_{\text{helpful}},\theta_{\text{harmless}},\theta_{\text{oasst}},\theta_{\text{webgpt}},\theta_{\text{summ.}}\}$ represent a preference model set. We utilize each preference model in the set to infer every data sample from the testset. We obtain the agreement testset which comprises data samples that agreed by the preference model set following Definition 1. The remaining samples form the disagreement testset. We analyze statistics and distribution of the reward difference (i.e., $r_{\theta}(x,y_{w})-r_{\theta}(x,y_{l})$) in reward values. It is worth noting that a positive reward difference $r_{\theta}(x,y_{w})-r_{\theta}(x,y_{l})>0$ indicates that the model accurately assigns a higher reward value to the winning sample.
Analysis of reward differences on each preference
Our findings indicate that the reward value distributions of different reward models are generally distinct. When examining the RM trained on different single preference datasets, we observe a significant diversity in reward values. In Figure LABEL:fig:reward_dist, we present the reward difference values of the first 50 test samples from the Harmless dataset as an example (Figure LABEL:fig:reward_a). Notably, even on the agreed samples, the baseline model provides different rewards. This observation suggests that the preference distribution learned from a single dataset by the reward model is highly diverse compared to others. Furthermore, Figure LABEL:fig:reward_b provides statistics of the reward difference across the entire test set, further supporting our statements regarding the diversity of reward values.
@@
Analysis of reward difference on training policies
MORE prevents the RM from outputting extreme reward values and maintains the reward accuracy (shown in Table 1). Specifically, we observe the reward values provided by RMs trained with different policies, as depicted in Figure 4. Compared with the sequential training scheme, MORE outputs reward values and variance with lower mean and variance. Connecting with our methodology, MORE mitigated the reward drifting. Hence, it tends to output a lower absolute reward signal as more accurate reward values according to ECE in Table 1. Furthermore, the hybrid training scheme also yields lower reward values compared to MORE. However, the reward accuracy is poorer in this case, which indicates that compressing the mean and variance of reward differences may degenerate the performance. These results suggest that the too-high or low reward differences may not be good in testing metrics. In contrast, MORE generates appropriate reward values, thereby enhancing the quality of reward modeling.
@@ 6 Discussion
Limitations
In this version, we have exclusively evaluated the MORE approach on the base model Pythia-1.4B. Further works on larger or smaller models are in progress. Besides, we randomly sample training samples from part of the dataset to avoid the effects of unbalanced datasets. The experiment of MORE handling the unbalancing mixed datasets will be presented in the next version. Throughout our experiments, we have found it necessary to utilize equation 8 in MORE, where we set $\alpha=0.5$ for all experiments. Further sensitive study on the value of setting $\alpha$ is in progress.
Future work
We have demonstrated the alignment ability of the reward model using the ECE metric (Zhu et al., 2023). However, the ultimate performance of the LLM, fine-tuned by the RM, remains a crucial aspect of our investigation. In our future work, we plan to complete the RL component in the RLHF pipeline. Additionally, we aim to answer several key questions regarding the reward model and MORE: 1) How do the number of diverse datasets and the scaling of base models impact the performance of MORE? 2) Does the RM trained by MORE yield improved results in RL tasks? 3) Can learning feedback from mixed diverse preference datasets enhance the performance of the LLM? Furthermore, we will delve deeper into understanding the implications of preference diversity in reward modeling. Gaining insights into the effects of preference diversity is crucial for designing algorithms that exhibit enhanced data efficiency.
Extension and Applications
The concept of learning shared preferences from mixed diverse preference datasets can be extended to DPO by incorporating the MORE loss (Rafailov et al., 2023). Furthermore, the approach of learning shared preferences through MORE can be implemented as a preference modeling pre-trained (PMP) paradigm (Askell et al., 2021). Subsequently, the RM trained using MORE can be easily fine-tuned to specific preferences (Cheng et al., 2023a). This flexibility allows for the adaptation of our approach to various applications and opens up new possibilities for preference modeling.
@@ 7 Related Work
RLHF has become the mainstream approach to align language models towards helpfulness, and harmlessness (Leike et al., 2018; Nakano et al., 2021b; Ouyang et al., 2022; Bai et al., 2022a). They all utilize an RM to align machine learning systems with human performance, which directly decides the performance of preference alignment. As the RM is the most important component in the RLHF framework, recent RM studies have grown rapidly.
The original goal of RM is to provide a scalar score to a model response and indicate the quality in equation 2, especially helpfulness and harmlessness. Due to the trade-off in quality aspects (Touvron et al., 2023a; Bai et al., 2022a), it can be challenging for a single RM to perform well in all aspects. Our work related to previous works handling multiple rewards and potential disagreement in preferences. For instance, Llama-2 (Touvron et al., 2023a) utilizes two separate RMs, one optimized for helpfulness and another for harmlessness. They mitigate the magnitude bias of the reward scalar with a margin loss, which provides a large margin for pairs with distinct responses, and a smaller one for those with similar responses. Multiple RMs can be utilized as majority voting or averaging (Jaques et al., 2020; Jang et al., 2023) in the PPO (Schulman et al., 2017). Wang et al. (2023a) introduces a Bayesian-based approach called d-PM to align language model with human preferences with disagreement. Cheng et al. (2023a) proposes to train a customized RM from the general RM to avoid disagreement from different preference domains. Furthermore, our theoretical intuition follows recent work DPO (Rafailov et al., 2023) and SLiC-HF (Zhao et al., 2023b) for preference alignment, which explores more straightforward methods to align language models with human preferences. Beyond the methodology, they have shown the RLHF framework is working as likelihood calibration tasks (Deng et al., 2020; Wang et al., 2023a; Azar et al., 2023), which proves that the reward values provided by the RM are also important.
@@ 8 Conclusion
This paper addresses the problem of diversified preferences in reward modeling for LLM alignment. Our experimental results demonstrate that RMs trained on different preference data sources exhibit substantial diversity in terms of reward distributions. This diversity derives from the intrinsic characteristics of the datasets, posing a challenge for RMs bridging the gaps among preference sources and better learning shared preference. To tackle this challenge, we propose a straightforward yet highly effective multi-objective reward model training scheme (MORE). Empirical results on five preference datasets shows that MORE can enhance the RM training to better capture and represent the shared preferences among multiple data sources. Furthermore, our proposed method outperforms conventional training schemes across three different evaluation metrics, highlighting its effectiveness and superiority.
Appendix A Decomposition of ranking loss
@@
Appendix B Experiments on Two Diverse Preference Datasets
@@
The results show that MORE implements comparable results with baseline methods in reward accuracy. However, the MORE is not the best on the ECE metric. This is because the number of diverse datasets is small (K=2), resulting in less diverse preference information. Hence, the reward drift mitigation task is solved inefficiently, making MORE degenerate to the vanilla method with rank loss as we demonstrated in Section A. Furthermore, we have achieved better results as shown in the main paper with $K=5$, indicating that MORE is efficient in handling diverse preferences in reward modeling.
Appendix C Diversity of Reward Values
In Figure LABEL:fig:additional-results, we provide additional experimental results as supplements of Figure 3.